OpenStudy (anonymous):

I have a sphere, x^2+y^2+z^2=25. I want to know the surface area above the cone described by z=sqrt(x^2+y^2). I have an idea of where to start, but...brain cramps. :P Notes to follow.

OpenStudy (anonymous):

First, I put the equation into cylindrical coordinates, substituting in the value of z provided for the cone.

OpenStudy (jhannybean):

interesting writing. hahaha

OpenStudy (anonymous):

:) I started out doing the Latex in Mathematica, but it was taking too long.

OpenStudy (anonymous):

Now I take the derivatives of r(u,v) with respect to u and v:

OpenStudy (anonymous):

Next the cross product. Maybe I should have done this in Mathematica! :)

OpenStudy (anonymous):

Only, should I have subtracted the cone?

OpenStudy (jhannybean):

wait a minute Your partial w.r.t u for the k component...

OpenStudy (jhannybean):

sin(u) becomes cos(u) when you take its partial wrt u, right?

OpenStudy (jhannybean):

You have 10sin(u) written there..

OpenStudy (anonymous):

Oopsie. If only we had a machine that could...oh wait. :)

OpenStudy (anonymous):

Aw man, I made even more biffs in in rv. Just a sec.

OpenStudy (anonymous):

Speedy. (!)

OpenStudy (jhannybean):

I was just correcting your error! And I don't think you can combine \(\sf \hat k\)

OpenStudy (jhannybean):

Well maybe you can if you factor out \(\sf 5\cos(u)\) but idk

OpenStudy (anonymous):

Wait a minute, shouldn't the first one be 5cos(u)cos(v) i ?

OpenStudy (jhannybean):

Oh yeah you're right,my bad.

OpenStudy (jhannybean):

I was focusedon the \(\sf \hat k\)

OpenStudy (anonymous):

No worries, I'm just happy I can even half-way participate in the conversation.

OpenStudy (jhannybean):

\[\sf \vec r_u = [5\cos(u)\cos(v)]\hat i +[5\cos(u)\sin(v)]\hat j + [5\cos(u)\cos(v) +5\cos(u)\sin(v)]\hat k\]

OpenStudy (anonymous):

Ok, I'm going to see if I get the same result on my partials, just a sec.

OpenStudy (jhannybean):

\[\sf \vec r(u,v) = [5\sin(u)\cos(v)]\hat i + [5\sin(u)\sin(v)]\hat j +[5\sin(u)\cos(v) +5\sin(u)\sin(v)]\hat k\]\[\sf \vec r_u = [5\cos(u)\cos(v)]\hat i +[5\cos(u)\sin(v)]\hat j + [5\cos(u)\cos(v) +5\cos(u)\sin(v)]\hat k\]\[\sf \vec r_v = [-5\sin(u)\sin(v)]\hat i + [5\sin(u)\sin(v)]\hat j +[5\sin(u)\sin(v)-5\sin(u)\cos(v)]\hat k\]

OpenStudy (jhannybean):

It kind of looked like you were trying to take second partials or something, deriving off of \(\sf \vec r_u\) or something.

OpenStudy (anonymous):

This is not territory for the feint of heart (or ADD, in my case). :P

OpenStudy (jhannybean):

i'm not exactly sure how to go about solving this problem either, but i Can help correct your steps! (Learning this as well)

OpenStudy (anonymous):

I appreciate the help, seriously. This semseter's been *brutal*.

OpenStudy (jhannybean):

And can you edit your question? Post what the exact question is/asking for?

OpenStudy (anonymous):

Sure.

OpenStudy (fibonaccichick666):

thanks

OpenStudy (anonymous):

Ok, here's the setup for my cross product:

OpenStudy (fibonaccichick666):

wait, can you post the actual question first, like word for word. I do not understand your question

OpenStudy (anonymous):

This is the question:

OpenStudy (jhannybean):

Oh, okay. this makes more sense

OpenStudy (jhannybean):

Did you find the intersection first?

OpenStudy (fibonaccichick666):

ohhh, makes alot more sense now

OpenStudy (jhannybean):

You have \[z=x^2 +y^2\] and \[z= \sqrt{25-x^2-y^2}\]

OpenStudy (jhannybean):

If you equate these two, you can find the intersection

OpenStudy (anonymous):

Ok, but that's just for the bounds, right?

OpenStudy (fibonaccichick666):

yep, bounds are important

OpenStudy (anonymous):

Yes, of course. What I mean is, I still need to parameterize, differentiate and take the cross product, right? I'd be happy to find I'm overcomplicating it, though.

OpenStudy (jhannybean):

I don't see why you'd be taking a cross product right away... but ofc i'd have to solve the problem in order to see why.

OpenStudy (anonymous):

What I'm looking for is the surface area of this:

OpenStudy (anonymous):

The orange part.

OpenStudy (anonymous):

Aw DERP. I waaay overcomplicated it!

OpenStudy (anonymous):

I can set it up as if I wanted the surface of the entire sphere, take the cross product and only integrate from the intersection of the cone on the bottom to the top of the sphere.

OpenStudy (jhannybean):

Whoop, idk what happened there. Yeah you did overcomplicate it.

OpenStudy (anonymous):

Ok, this shouldn't take much longer now. Wow, little misunderstandings can take up a TON of time. :P

OpenStudy (anonymous):

ON the bright side, I guess I won't make that mistake again.

OpenStudy (anonymous):

Ok, so here goes:

OpenStudy (fibonaccichick666):

lol yea, you only need the intersection points :) then you can find the central angle

OpenStudy (fibonaccichick666):

then you can find the surface area based off of the area of the circle, could potentially do it without calculus

OpenStudy (anonymous):

Yeah, but when the final gets here in two week, I'd like to present it in the way the prof want to see it. :)

OpenStudy (fibonaccichick666):

well... yea... you can do that using a surface integral

OpenStudy (fibonaccichick666):

but you should get two identical answers

OpenStudy (anonymous):

Working through it as we speak.

OpenStudy (fibonaccichick666):

okee dokey

OpenStudy (fibonaccichick666):

( you can use the easy geometry way to check your answer though )

OpenStudy (anonymous):

Yup.

OpenStudy (camper4834):

im sorry but can you explain why you did what you did calc3 student

OpenStudy (camper4834):

instructor never explained finding surface area

OpenStudy (anonymous):

Well, if I could do that, I'd be done. :) I will as soon as I get through this round, though. Our class didn't do very well on it either. If you want a great discussion of it, though, Khan Acadamy goes into pretty good detail: https://www.khanacademy.org/math/multivariable-calculus/surface-integrals

OpenStudy (fibonaccichick666):

@camper4834, check this out http://tutorial.math.lamar.edu/Classes/CalcIII/SurfaceArea.aspx

OpenStudy (fibonaccichick666):

yep, those are my top two choices to explain just about everything calculus

OpenStudy (anonymous):

The Paul's link is great, too.

OpenStudy (fibonaccichick666):

the only unfortunate thing is it doesn't explain the derivation

OpenStudy (anonymous):

For how much gets left out of books and not explained in class, it's a wonder anyone learns anything at all.

OpenStudy (fibonaccichick666):

true. It gets worse the higher you go

OpenStudy (anonymous):

That's encouraging. On the bright side (pun intended), the full moon is just coming up here, and it's *beautiful*. And so are my two daughters, who I'm really starting to miss. I'm going to take a break for dinner, and I'll be back to finish this. Thanks, everyone for your help so far.

OpenStudy (fibonaccichick666):

But, I hope not to be too lacking when I teach

OpenStudy (fibonaccichick666):

lol, and yea, calc III sucks for memorization. There are some practice tests you can find online and use

OpenStudy (anonymous):

I hear you @FibonacciChick666. I've never wanted to start teaching again so much as since I've gone back to school. It's like I feel compelled to save the world. And that's funny...the link you gave me is from my school--I'm at the University of Utah.

OpenStudy (fibonaccichick666):

I spent 72 hours the week of the final teaching myself the information. I had a very bad graduate student for the class. lol first link that came up, but then if you can do those, you should be fine. :)

OpenStudy (fibonaccichick666):

and yea, bad teachers make me want to teach, but common core man.... it is a disgrace

OpenStudy (fibonaccichick666):

I am still not sure how I can teach effectively with those limitations

OpenStudy (anonymous):

I'll never understand the bureaucratic hostility towards an iterative approach to understanding. It works *great* on Khan. Yet my physics dept continues to lose 85% of their declared majors by teaching the topics same way Newton would have learned them (had he not invented them).

OpenStudy (fibonaccichick666):

.... yep. But don't worry hs graduates won't be able to do algebra 1 in a few years.

OpenStudy (fibonaccichick666):

let alone physics

OpenStudy (anonymous):

Perhaps the problem is precisely that people who are great teachers can't stand the thought of doing it wrong. But I digress. Here's what I have so far:

OpenStudy (fibonaccichick666):

meh maybe, but anywho, shouldn't your phi(since you bisected the chord) be twice what you got? Argue it if you disagree please

OpenStudy (fibonaccichick666):

also, what program are you using????? I quite like it

OpenStudy (anonymous):

The program is Photoshop. I've been using it for ...well since version 2.0. I'm what you'd call a non-traditional student (in lots of ways). As for the chord bit, I've no idea.

OpenStudy (fibonaccichick666):

lol I didn't know photoshop could do that. anyways here is what I am seeing I assume this is your picture, if it is correct? 

OpenStudy (fibonaccichick666):

so then your representative looks like this, but what you have is the second